OpenStudy (anonymous):

Find the 40th term of the sequence 2, -12, 72,.... Write the arithmetic sequence that has 4 means between 5 and 15

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

Find the 40th term of the sequence What kind of sequence? arithmatic or geometric sequence?

OpenStudy (anonymous):

arithmitic

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Arithmetic_progression Use the very first formula 4oth term is 2+(40-1)*((-12)-2)=-544 but i think it geo something dont seem right.

OpenStudy (anonymous):

it actualy doesnt say..i actually assumed..but if it doesnt seem right then it is geo

OpenStudy (anonymous):

well geo seems fishy too go with -544 :P I am so sorry if it is wrong:(

OpenStudy (anonymous):

ok..thanks! how about #2?

OpenStudy (anonymous):

i am sorry got no idea on second one i am googling tho :P

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

http://www.regentsprep.org/Regents/math/algtrig/ATP2/ArithSeq.htm Look at problem 9 :D Write the arithmetic sequence that has 4 means between 5 and 15 5, , , , 15

OpenStudy (anonymous):

15=5+(6-1)*d sfd d=2 now insert terms using d by adding 2 to the 5 that be 2nd term and you get 7 then add 2 to 7 to get 9 that be your 3rd and so on and so on :) 5, 7 , 9,11 ,13, 15 ^ ^ ^ ^

OpenStudy (anonymous):

thanks alot! :)